Harmonic Mean Example
Investor buys $3,000 of a stock at the end of Month 1 at $20 a
share, and $3,000 at the end of Month 2 at $25 per share.
What is the average cost per share of stock?
Answer #1

General method:

At the end of first month, no of stocks purchased = 3000/20 = 150

At the end of second month, no of stocks purchased = 3000/25 = 120.

total amount invested = 6000, total share purchased = 150+120 = 270.

so, average cost per share = 6000/270 = 200/9 = 2.22 $ per share.

Using harmonic mean:

Harmonic mean of a & b = 2ab/(a+b)

So, since amount invested is same, avg cost = 2•20•25/(20+25) = 1000/45 = 200/9 = 22.22